OpenStudy (anonymous):

MEDAL!! Select the equation of the line that passes through the point (–2, 1) and has slope 3 in point-slope form. A. (x + 2) = 3(y – 1) B. (y + 1) = 3(x – 2) C. (y – 1) = 3(x + 2) D. (x – 2) = 3(y + 1)

OpenStudy (anonymous):

IMPORTANT LINE RELATED EQUATIONS TO KNOW AND MEMORIZE slope formula m= slope/ gradiant -- same thing \[m=\frac{y_2-y_1}{x_2-x_1}\] standard formula \[Ax+By=C\] point-slope formula \[y-y_1=m(x-x_1)\] slope-intercept formula b= y-intercept -- in the form of (0,y) \[y=mx+b\]

OpenStudy (anonymous):

ik! its just im still not getting the right answer

OpenStudy (anonymous):

which answer do you think it is?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

look carefully at the point slope form i have written above are the x and y in the proper locations?

OpenStudy (anonymous):

b!!!!

OpenStudy (anonymous):

x - (-2)=?

OpenStudy (anonymous):

y-1 =?

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

yea, although i get the feeling you were just guessing o well

OpenStudy (anonymous):

Noooo reallly thought it wAS D!!! Then my sister came over and she thought it was b. But then i thought it was c.